Do-Well bonds have a face value of $1,000 and are currently quoted at 86.725. The bonds have a coupon rate of 6.5 percent. What is the current yield on these bonds?
Answer #1

Price = 86.725% of 1000 = 867.25

Coupon = 0.065 * 1000 = 65

Current yield = (Annual coupon / price) * 100

Current yield = (65 / 867.25) * 100

Current yield = 7.50%